OpenStudy (anonymous):

Which is equivalent to (-2 + 4i)(7 + 2i)?

OpenStudy (anonymous):

the choice that has the correct one....

OpenStudy (anonymous):

AINT NO CHOICE

OpenStudy (anonymous):

9+6i

OpenStudy (paxpolaris):

\[ (-2 + 4i)(7 + 2i)=-14-8-4i+28i=\Large-22+24i\]

OpenStudy (anonymous):

that one....